# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Coding

### v1

@@ 5.4.3. The creators whose works make up the training data
The next largest group, 156 respondents (28.2%), believed that rights in the generated code should belong to the individuals who created the code on which the GenAI model was trained. A common theme among such responses was a sense of fairness or justice, given GenAI’s reliance on training data. As $R_{77}$ put it, “[I]f the code is extremely similar or identical to the training data[,] then the copyright should belong to the respective creator of this training data.” $R_{13}$ wrote, “[I]f these humans had not written this code, this model would not exist,” while $R_{25}$ echoed, “Without training data, AI can not do anything.” Some responses expressed this view in stronger language, with $R_{31}$, for example, calling large language models “plagiarism engines.”
Some respondents, by contrast, invoked the collaborative nature of development as a reason to deny copyright ownership claims by developers of the code used in the training data. $R_{157}$ noted, “[T]he neural networks require so much input for them to work that none of the input is really deterministic to the end work. […] So the copyright can not belong to the people that just provided the training data. [If that were true], the copyright of [a] book could belong also to the authors of all the books that the author had read before.” $R_{157}$ continued by providing another analogy from the creative arts: “[We don’t give] copyright to the people that cleaned the theater and people that create the music[al] instruments and so on.” Respondents also identified practical problems with this ownership claim. $R_{187}$ said, “AI models are trained using [a] large [volume] of source code (some of which may be extremely similar) from different sources, so it is difficult to say which parts of the training material the generated code belongs to.” In other words, an exact mapping from the training examples to the generated output might be difficult, if not impossible, to accomplish, particularly in situations with many similar or identical code instances (such as GitHub forks). As $R_{502}$ asked, “[H]ow can we be so sure whether that code was originally written by that user?”
Finding 20: Some developers believed that fairness requires giving ownership of generated code to those developers who created the training data. However, others invoked the collaborative nature of development as a reason to deny copyright claims by those whose code was used in the training data and highlighted the practical problems with mapping training data to generated code in order to determine ownership.
@@ 5.4.4. The creators of the model
The least popular of the provided options, ranked even below the “don’t know/no opinion” option, was that the creators of GenAI models should own the output, with only 49 participants (8.8%) indicating that they believed this to be the case. To justify this position, $R_{1}$ cited the “resources, time, money, effort and research that goes into creating novel systems.” Similarly, $R_{441}$ believed that “the creators of the model should receive monetary compensation for training expenses.”
By contrast, other respondents, such as $R_{306}$, indicated that while the ownership question is complicated, model creators were the one group with no real claim: “Again, it is difficult to be sure. What I do believe, however, is that it should never belong to the creators of the model, as they have no say on the output of the model, only its inner workings.” $R_{98}$ expressed a similar sentiment: “The creators of the model are definitely not the copyright owners—LLMs are just text generators and as such can technically generate any text. You can force [a] LLM to ‘write’ literally any sequence of words you want—that [output] can’t be copyrighted [by the model].” $R_{133}$ said “I can say very confidently that model creators should NOT get attribution for anything their model creates since they only aggregate the data into the model.”
@@ 5.4.5. Other ownership claims
A few respondents chose multiple options because they believed that copyright ownership should be distributed among multiple groups. For example, $R_{191}$ called generated code “a co[-]creation where we would have a percentage of ownership to apply. […] The prompt, however, [would] belong entirely to [the] user.” Similarly, $R_{264}$ pointed out that “Without original code, we couldn’t train the AI tool; without the creator of the model, we will never have such a tool; without the [prompter], we couldn’t get an ideal result. Everyone contributes to the AI generated results, so I think everyone should have some copyrights.” $R_{527}$ suggested, “I would opt [for] […] a shared benefits model, like [with] royalties in the entertainment industry, where the actual code owner gets a fair share, the model owner, and the end user who was capable of adding the right prompt into the model […].” Such views were uncommon, and no respondent proposed a system for determining an appropriate allocation of ownership rights or the implications thereof.
@@
Interestingly, a handful of respondents ($R_{86}$, $R_{489}$, $R_{593}$, $R_{715}$) advocated that the model itself should maintain rights to the content it generates, at least under certain conditions.
In the end, as $R_{54}$ noted, as GenAI becomes an increasingly common feature of all software, the question of who owns GenAI output may cease to matter: “[T]he issue [of ownership] becomes more clouded, though, when you consider the current situation, where most generative AI is done through a service[, n]ot directly with the model. Therefore the service that is providing the generated code or response can define in their own terms [of service] whether they own [the output] or not.”
@@ 5.5. On the ownership and copying of prompts
Much of the current GenAI litigation, as of this writing, relates to the unauthorized use of preexisting work in training data or the copyright status of generated content. We were, however, also interested in understanding developers’ views on the ownership of prompts, which, depending on their complexity, could be protected by copyright law. (As noted above, however, platforms’ Terms of Service might govern users’ rights in this regard.) About half of the 554 respondents familiar with GenAI tools (278, 50.2%) were indifferent as to whether a prompt they developed could be used by someone else without their permission, with 88 respondents (15.9%) indicating that they would be pleased if such use occurred and 63 (11.4%) extremely pleased. An additional 89 respondents (16.1%) would be displeased by the use of their prompts, and 36 (6.5%) extremely displeased. Overall, as shown in Figure 11, most developers in our survey stated that they would not care (or would even be pleased) if their prompts were reused by others, even without permission.
Finding 23: Most respondents (77.4%) indicated that they would be indifferent or even pleased if their prompts were copied and used by others, even without their permission.
@@ 5.6. Perceived risk of potential copyright infringement
Ultimately, many developers perceived that the likelihood of legal enforcement resulting from improper usage of GenAI was low. $RI_{149}$ stated their belief: “Nobody’s really looking, and if [copyrighted code] managed to get into a training data set somewhere, then [the copyright owner is] to blame.” $RI_{39}$ stated in an interview: “If I technically am violating somebody’s copyright for code that is never going to be pushed anywhere public, it’s very implausible that I would be sued for that.” $RI_{31}$ noted that they belonged to “a fairly small organization. We’re not high on the radar. I don’t think that we’re at any legal risk of anything. If anybody’s going to get sued, it’s not going to be us. It’s going to be other people.” $RI_{149}$ expressed a similar view: “[T]he likelihood of detection is so small that it’s no longer a key concern.”
@@
Finding 24: Individuals and smaller organizations may perceive less practical risk from using code from GenAI due to the low likelihood of any code that potentially infringes on another’s copyright or violates another’s license being discovered.
Table 6. Main findings for RQ2: Developer perceptions of copyright issues related to GenAI
5.1 Thoughts on Litigation
F11
While about half of participants anticipated that litigation would result in limitations on the use of
GenAI, a vocal minority suggested that imposing such limitations would be difficult if not impossible,
and others urged restraint, given the pace of technological developments.
5.2 Using one’s own code in models’ training data or as output
5.2.1 Sentiments regarding the inclusion of one’s own code in training data
F12
Developers’ views on having their code included in models’ training data was context-dependent.
If the code was open source, many developers were indifferent or even pleased to have their work
included; this was not true for proprietary code. Developers also expressed concerns about the quality
of code incorporated into the training data and whether increased use of GenAI tools would disincentivize
developers from contribution to the OSS community.
5.2.2 Conditions around use of one’s code in training data
F13
Some respondents suggested that training data contributors should recieve monetary compensation or
other benefits, such as free access to the trained model or free credits for the platform. However, forcing
AI companies to provide monetary compensation to all contributors of the training data, while seemingly
fair, would place an additional hurdle potentially blocking smaller AI companies from entering the market.
F14
Respondents generally wanted licensing terms to be complied with by GenAI models, especially when
producing code kidential to or very similar to work in the training set.
5.2.3 Fair use and replicated code
F15
Some developers believed that the use of code as training data should be considered fair use by default and
that the onus should be on developers to protect code that they do not wish to be included in training data.
5.3 AI-generated output that resembles existing code
F16
A group of developers did not see how GenAI technology was doing anything significantly different from
traditional means of creating code, which reuses existing code created by others. The only distinction is the
scale and ease at which it can be done.
F17
Some developers’ views on the copying of code track U.S. copyright law: basic functions and building blocks
are not protectable and should be free for all to use.
5.4 On the ownership of AI-generated code
5.4.1 No one-works are in the public domain
F18
The largest group of developers (242, 43.68%) thought that the output of GenAI belongs in the public domain.
Some believe this because it is consistent with the collaborative nature of open-source software development,
while others believe that, because AI models are not human, the output of those models cannot be subject to
copyright law.
5.4.2 The prompting developer or their employer
F19
The second largerst group of developers (201, 36.3%) believed that a user of a GenAI tool should own the
copyright to code generated with that tool because it is the user’s contriubtions—creating the prompt and,
typically, editing and further refining the result—that create the code. In other words, GenAI was seen as
nothing more than a new tool for creating content.
5.4.3 The creators whose works make up the training data
F20
Some developers believed that fairness requires giving ownership of generated code to those developers who
created the training data. However, others invoked the collaborative nature of development as a reason to deny
copyright claims by those whose code was used in the training data and highlighted the practical problems
with mapping training data to generated code in order to determine ownership.
5.4.4 The creators of the model
F21
The model vendors who developed GenAI models were the one group that the fewest participants (49)
believed had an ownership claim to generated outputs.
5.4.5 Other ownership claims
F22
Some respondents believed that ownership depended on various factors, including the relevant jurisdiction,
the amount of code generated, and how much the prompting developer contributed to the result. Others
believed that multiple parties share claims to ownership and that a royalty-style system should be put into place.
5.5 On the ownership and copying of prompts
F23
Most respondents (77.4%) indicated that they would be indifferent or even pleased if their prompts were copied
and used by others, even without their permission.
5.6 Perceived risk of potential copyright infringement
F24
Individuals and smaller organizations may perceive less practical risk from using code from GenAI due to the low
likelihood of any code that potentially infringes on another’s copyright or violates another’s license being discovered.
@@ 6. RQ3: Other legal concerns that developers anticipate as the use of generative AI increases
RQ3 aimed to discover additional legal concerns (other than those relating to copyright) that developers anticipate as the use of GenAI increases. The possibility of liability may already be motivating organizations; $R_{34}$ indicated, “[My organization is] accounting for the possibility of legal risks and [has] an open source alternative if [O]penAI or [M]icrosoft are sued.” Here we report these additional concerns that were on the minds of the surveyed developers. These include the aforementioned question of liability, the generation of non-code content, the creation of malicious content, risks involving data privacy, and the use of GenAI for software testing activities. A summary of key findings can be found in Table 7.
@@ 6.1. Liability concerns
Respondents mentioned concerns about possible tort liability related to the use of GenAI. Product liability was mentioned by 14 respondents (2.5%), and liability for code generated with bugs or defects was mentioned by 13 respondents (2.4%). Developers were concerned about who would be held accountable for security issues or other bugs introduced by AI-generated code. $R_{0}$ expressed that they “could see in the near future running [into] issues with security breaches caused by an AI code generator making sub-optimal choices. A contrived example[,] but imagine someone uses the AI to build a log in. The AI suggests that you store passwords in plain text in your database which is later leaked. The developer could argue the model creator (e.g. [O]penAI) was responsible for the breach.”
Other respondents identified potential issues with AI agents, such as consumer-facing chatbots that purport to engage in contractual agreements with users (Belanger, 2024; Malley, 2023) or AI agents that appear to operate autonomously. $R_{149}$ said that “[f]unction calling”—using LLMs “to write and call functions”—was a potential concern (see examples in (AutoGPT, [n. d.]; gptengineer, [n. d.])) and asked, “[W]hat happens if those called (which may have real world effects) cause harm—who is responsible?” $R_{204}$ raised the similar concern of rogue agents, which could be given a set of directives to autonomously carry out for malicious or nefarious purposes.
Many respondents explicitly mentioned that the use of generative technologies in important systems and domains posed a particularly high risk. $R_{109}$ asked “[i]f any medical, aviation, security or defen[s]e-based grade code is written by such systems, we better double check the work, and who takes the fall when such a system reaches production and fails?” $R_{245}$ commented more broadly: “Software liability is complicated, and although right now there seems to be a kind of stable equilibrium in terms of how liability is determined, AI may upend that equilibrium if we see individuals and companies with very large code bases that are mostly not authored by human beings. Some of these parties will inevitably end up trying to make the claim that liability should not attach in circumstances where errors such as ‘negligence’ were actually never carried [out] by a human being. This strikes me as a major concern for some types of software used in aerospace or medicine, etc.”
Finally, 13 respondents (2.35%) were concerned that code might be generated without sufficient consideration of compliance issues, such as “ignoring GDPR [the EU’s General Data Protection Regulation]” ($R_{190}$). $R_{665}$ was concerned that “[i]f an AI generates specifications for a project, it may not consider existing laws and [so] generate software which does not comply with regulations.” Several respondents raised the related concern of using GenAI to draft legal documents such as software licenses, terms of service, and privacy policies, which could not only lead to an undesirable proliferation of such documents but also may result in language that does not have the desired legal effect. Similar problems of using AI to interpret or clarify legal standards were identified by $R_{527}$.
@@
6.2. Generating content other than code
Although our survey and interview questions focused primarily on copyright issues relating to code, 92 respondents (16.6%) also noted concerns about copyright issues relating to images and other types of content generated by AI, which might arise for developers in the context of “logo design [and] art generation for video games” ($R_{232}$) or fonts, icons, and images used in front-end web development, to take a few examples. (The legal status of outputs of image generation through AI tools such as MidJourney (Midjourney, [n. d.]), StableDiffusion (Stability AI, [n. d.]), and DALL-E (DALL-E 3, [n. d.]) is currently a subject of debate (get, 2023; Andersen v. Stability AI Ltd., [n. d.]).)
In a similar vein, 23 respondents (4.2%) noted that tasks related to the development of user interfaces (UI) could lead to legal issues. $R_{54}$ asserted that “SVG images and logos are often created and defined directly with the code. Likewise design elements, CSS, and theme files are also involved in software design but are much easier targets for legal copyright as visual similarities are easier to point out.” $R_{30}$ added an additional concern: “I think relying on generative AI for UI/UX design can easily lead to essentially identical works.”
Finding 26: The generation of other types of media (images, fonts, videos, audio, etc.) and theme/style-sheets can also pose potential legal challenges for development teams since such assets are often used in product creation.
@@ 6.3. Generation of malicious content
Seventeen survey respondents also highlighted the potential to use GenAI tools to produce a variety of malicous content, including malware or viruses, disinformation, illicit or unsavory material, and propaganda. While not a novel insight (Fritsch et al., 2022; Goldstein et al., 2024; Pa Pa et al., 2023), developers are aware that the technical and skill barriers that have existed in creating malicious content have been lowered, enabling script-kiddies to carry out more sophisticated attacks with little to no domain knowledge. $R_{489}$ suggested that “AI [could be] used to maliciously attack someone (constant spam, emails, etc. too fast to fight back),” and $R_{411}$ mentioned that generative AI technologies could be used for carrying out search engine optimization (SEO) hacks. $R_{144}$ also raised the concern that models could generate malicious or toxic content that the end user did not solicit, which “could lead to legal repercussions in many EU countries.” While many mainstream LLMs have guardrails in place to prevent the generation of this type of content (OpenAI, 2024), unrestricted, open-weight models such as the Dolphin family of models (Sarkari, 2024) and alternatives such as WormGPT (Newsroom, 2023) can still be used by malicious actors, and jailbreaking techniques can be used to bypass safety rails even for more widely used models (Wei et al., 2024; Liu et al., 2024).
Finding 27: Respondents identified the purposeful generation of malicious content as a societal problem and legal concern going forward.
@@ 6.4. Data privacy, information leakage, and discovery of system vulnerabilities
Information leaks were mentioned by 33 (6%) respondents, and data privacy concerns by an additional 15 (2.7%). This concern has been explored in the existing literature (Abascal et al., 2023; Carlini et al., 2020, 2021), but it is of note that it is also a real world consideration for developers when choosing which GenAI tools to trust and use. Possible issues included proprietary code or information being produced by generative AI, leaked secrets such as API tokens and passwords, and the release of personally identifiable information. This risk results, as $R_{447}$ put it, from the fact that the “devs […] didn’t realize [sensitive information would] get hoovered up” in the training data. $R_{501}$ echoed this sentiment, noting that models were “[p]icking up private material that was committed accidentally.”
Finding 28: Developers and other stakeholders did not intend for or anticipate that their data would be collected en masse, so developers are concerned that private and sensitive information can easily make its way into training sets.
Similarly, there were concerns that generative models could be used by malicious actors for hacking (6), reverse engineering (6), and jailbreaking (1) systems. $R_{393}$ said that people could use “AI to hack/discover security vulnerabilities […].” $R_{139}$ said that people might gain “access to secure systems through intelligent vulnerability discovery and exploitation.” $R_{352}$ worried that AI tooling “could make it easier to jailbreak secure applications by learning salts and similar common practices in infrastructure to breach a site or app.”
Finding 29: Generative AI, particularly LLMs, can not only leak information that was contained in the original training data but can also be used by malicious actors to exploit systems.
Respondents were also cognizant of the dangers of including private or proprietary information in prompts, fearing that those same prompts would be used in training sets of future models. $RI_{31}$ explained, “Because I’m almost only doing work stuff, and in the enterprise environment, I can’t run the risk of the training set pulling security keys and that kind of thing. Because it’s been demonstrated that you can trick those things into outputting data from the training set.” $RI_{516}$ said, similarly, that “I use [GenAI] only for things that are very generic and do not use code that would not already go in the open.”

### v3

@@ 5.4.3. The creators whose works make up the training data
The next largest group, 156 respondents (28%), believed that rights in the generated code should belong to the individuals who created the code on which the GenAI model was trained. A common theme among such responses was a sense of fairness or justice, given GenAI’s reliance on training data. As $R_{77}$ put it, “[I]f the code is extremely similar or identical to the training data[,] then the copyright should belong to the respective creator of this training data.” $R_{13}$ wrote, “[I]f these humans had not written this code, this model would not exist.”
Some respondents, by contrast, invoked the collaborative nature of development as a reason to deny copyright ownership claims by developers of the code used in the training data. $R_{157}$ noted, “[T]he neural networks require so much input for them to work that none of the input is really deterministic to the end work. […] So the copyright can not belong to the people that just provided the training data. [If that were true], the copyright of [a] book could belong also to the authors of all the books that the author had read before.” $R_{157}$ continued by providing another analogy from the creative arts: “[We don’t give] copyright to the people that cleaned the theater and people that create the music[al] instruments and so on.”
Respondents also noted that it may be difficult, if not impossible, to extract an exact mapping from the training examples to the generated output, particularly in situations with many similar or identical code instances (such as GitHub forks). As $R_{502}$ asked, “[H]ow can we be so sure whether that code was originally written by that user?”
Finding 20: Some developers believed that fairness requires giving ownership of generated code to those developers who created the training data. However, others invoked the collaborative nature of development to deny copyright claims by those whose code was used in the training data and highlighted the practical problems with mapping training data to generated code to determine ownership.
@@ 5.4.4. The creators of the model
The least popular of the provided options, ranked even below the “don’t know/no opinion” option, was that the creators of GenAI models should own the output, with only 49 participants (9%) indicating that they believed this to be the case. $R_{441}$ believed that “the creators of the model should receive monetary compensation for training expenses.”
By contrast, other respondents, such as $R_{306}$, indicated that while the ownership question is complicated, model creators were the one group with no real claim: “Again, it is difficult to be sure. What I do believe, however, is that it should never belong to the creators of the model, as they have no say on the output of the model, only its inner workings.” $R_{98}$ expressed a similar sentiment: “The creators of the model are definitely not the copyright owners—LLMs are just text generators and as such can technically generate any text. You can force [a] LLM to ‘write’ literally any sequence of words you want—that [output] can’t be copyrighted [by the model].”
@@ 5.4.5. Other ownership claims
A few respondents chose multiple options because they believed that copyright ownership should be distributed among multiple groups. For example, $R_{191}$ called generated code “a co[-]creation where we would have a percentage of ownership to apply. […] The prompt, however, [would] belong entirely to [the] user.” $R_{527}$ suggested, “I would opt [for] […] a shared benefits model, like [with] royalties in the entertainment industry, where the actual code owner gets a fair share, the model owner, and the end user who was capable of adding the right prompt into the model […].” Such views were uncommon, and no respondent proposed a system for determining an appropriate allocation of ownership rights or the implications thereof.
@@
In the end, as $R_{54}$ noted, as GenAI becomes an increasingly common feature of all software, the question of who owns GenAI output may cease to matter: “[T]he issue [of ownership] becomes more clouded, though, when you consider the current situation, where most generative AI is done through a service[, n]ot directly with the model. Therefore, the service that is providing the generated code or response can define in their own terms [of service] whether they own [the output] or not.”
@@ 5.5. On the ownership and copying of prompts
Much of the current GenAI litigation, as of this writing, relates to the unauthorized use of preexisting work in training data or the copyright status of generated content. We were, however, also interested in understanding developers’ views on the ownership of prompts, which, depending on their complexity, could be protected by copyright law. (As noted above, however, platforms’ Terms of Service might govern users’ rights in this regard.) About half of the 554 respondents familiar with GenAI tools (278, 50%) were indifferent as to whether a prompt they developed could be used by someone else without their permission, with 88 respondents (16%) indicating that they would be pleased if such use occurred and 63 (11%) extremely pleased. An additional 89 respondents (16%) would be displeased by the use of their prompts, and 36 (7%) extremely displeased. Overall, as shown in Figure 8, most developers in our survey stated that they would not care (or would even be pleased) if their prompts were reused by others, even without permission.
Finding 23: Most respondents (77%) indicated that they would be indifferent or even pleased if their prompts were copied and used by others, even without their permission.
@@ 5.6. Perceived risk of potential copyright infringement
Ultimately, in our follow-up interviews, five of seven respondents perceived that the likelihood of legal enforcement resulting from improper usage of GenAI was low. $RI_{39}$ stated in an interview: “If I technically am violating somebody’s copyright for code that is never going to be pushed anywhere public, it’s very implausible that I would be sued for that.” $RI_{31}$ noted that they belonged to “a fairly small organization. We’re not high on the radar. I don’t think that we’re at any legal risk of anything. If anybody’s going to get sued, it’s not going to be us. It’s going to be other people.” $RI_{149}$ expressed a similar view: “[T]he likelihood of detection is so small that it’s no longer a key concern.”
@@
Finding 24: Individuals and smaller organizations may perceive limited practical legal risk from using AI-generated code due to the low likelihood that any code that potentially infringes on another’s copyright or violates another’s license would be discovered.
@@ 6. RQ3: Other legal concerns that developers anticipate as the use of generative AI increases
The legal issues other than those related to copyright law that GenAI raises are apparent. RQ3 aimed to discover which of these concerns were top-of-mind among the developers in our study. Developers identified a broad range of issues, including those relating to tort liability, malicious use, and data privacy concerns. Findings 25-29 answer this research question, a summary of which is also available in our replication package (ano, 2024).
@@ 6.1. Liability concerns
Respondents mentioned concerns about possible tort liability (legal responsibility for harmful actions) related to the use of GenAI. Product liability (the claim, for example, that a product is defective or harmful) was mentioned by 14 respondents (3%), and liability for code generated with bugs or defects was mentioned by 13 respondents (2%). In particular, developers were concerned about who would be held accountable for security issues or other bugs introduced by AI-generated code. $R_{0}$ expressed that they “could see in the near future running [into] issues with security breaches caused by an AI code generator making sub-optimal choices. A contrived example[,] but imagine someone uses the AI to build a log-in. The AI suggests that you store passwords in plain text in your database which is later leaked. The developer could argue the model creator (e.g. [O]penAI) was responsible for the breach.”
Other respondents identified potential issues with AI agents, such as consumer-facing chatbots that purport to engage in contractual agreements with users (Belanger, 2024; Malley, 2023) or AI agents that appear to operate autonomously, such that parties may attempt to disclaim responsibility for harmful effects. $R_{149}$ said that “[f]unction calling”—using LLMs “to write and call functions”—was a potential concern (see examples in (AutoGPT, [n. d.]; gptengineer, [n. d.])) and asked, “[W]hat happens if those called (which may have real world effects) cause harm—who is responsible?” Respondents also highlighted the risk posed by using GenAI in safety/mission-critical systems and domains, such as aviation, security, defense, and medical fields.
Finally, 13 respondents (2%) were concerned that code might be generated without sufficient consideration of compliance issues, such as “ignoring GDPR [the EU’s General Data Protection Regulation]” ($R_{190}$). $R_{665}$ was concerned that “[i]f an AI generates specifications for a project, it may not consider existing laws and [so] generate software which does not comply with regulations.” Several respondents raised the related concern of using GenAI to draft legal documents such as software licenses, terms of service, and privacy policies, which could not only lead to an undesirable proliferation of such documents but, if not subject to sufficient human review, also may result in language that does not have the desired legal effect.
@@ 6.2. Generation of malicious content
Seventeen survey respondents also highlighted the potential to use GenAI tools to produce a variety of malicious content, including malware or viruses, disinformation, illicit or unsavory material, and propaganda. While not a novel insight (Fritsch et al., 2022; Goldstein et al., 2024; Pa Pa et al., 2023), developers are aware that pre-existing technical and skill barriers to creating malicious content have been lowered, enabling less experienced individuals to carry out more sophisticated attacks with little to no domain knowledge. While many mainstream LLMs have guardrails in place to prevent the generation of this type of content (OpenAI, 2024), unrestricted, open-weight models such as the Dolphin family of models (Sarkari, 2024) and alternatives such as WormGPT (Newsroom, 2023) can still be used by malicious actors, and jailbreaking techniques can be used to bypass safety rails even for more widely used models (Wei et al., 2024; Liu et al., 2024).
Finding 26: Respondents identified the purposeful generation of malicious content as a societal problem and legal concern going forward.
@@ 6.3. Data privacy, information leakage, and discovery of system vulnerabilities
Information leaks were mentioned by 33 (6%) respondents, and data privacy concerns by an additional 15 (3%). This concern has been explored in the existing literature (Abascal et al., 2023; Carlini et al., 2020, 2021), but it is of note that it is also a real-world consideration for developers when choosing which GenAI tools to trust and use. Possible issues include proprietary code or information being produced by generative AI, leaked secrets such as API tokens and passwords, and the release of personally identifiable information. This risk results, as $R_{447}$ put it, from the fact that the “devs […] didn’t realize [sensitive information would] get hoovered up” in the training data.
Finding 27: Developers and other stakeholders did not intend for or anticipate that their data would be collected en masse, so developers are concerned that private and sensitive information can easily make its way into training sets.
Similarly, there were concerns that generative models could be used by malicious actors for hacking (6), reverse engineering (6), and jailbreaking (1) systems. $R_{393}$ said that people could use “AI to hack/discover security vulnerabilities […].”
Finding 28: Generative AI, particularly LLMs, can not only leak information that was contained in the original training data but can also be used by malicious actors to exploit systems.
Respondents were also cognizant of the dangers of including private or proprietary information in prompts, fearing that those same prompts would be used in training sets of future models. $RI_{31}$ explained, “Because I’m almost only doing work stuff, and in the enterprise environment, I can’t run the risk of the training set pulling security keys and that kind of thing. Because it’s been demonstrated that you can trick those things into outputting data from the training set.”